# Unified Unbiased Variance Estimation for MMD: Robust Finite-Sample Performance with Imbalanced Data and Exact Acceleration under Null and Alternative Hypotheses

### v1

@@ 4.2 Accelerated computation of variance components
$\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}$ $=\sum_{i=1}^{n}(R_{i}^{\mathbf{X}}+L_{i}^{\mathbf{X}})^{2},$ (60)
@@
$\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}$ $=\sum_{j=1}^{m}(R_{j}^{\mathbf{Y}}+L_{j}^{\mathbf{Y}})^{2},$ (63)
@@
$\|\mathbf{K}_{XY}\|_{F}^{2}$ $=\sum_{i=1}^{n}(\mathcal{A}_{x\to y}(i)+\mathcal{Z}_{x\to y}(i))^{2},$ (66)
@@
Based on the theoretical analysis, Algorithm 4 computes the second-order kernel quantities from prefix–suffix row sums, while Algorithm 5 presents an accelerated procedure for computing the Maximum Mean Discrepancy (MMD) with the Laplace kernel together with its full Hoeffding variance estimator, reducing the computational complexity from the naive $\mathcal{O}(N^{2})$ to $\mathcal{O}(n\log n+m\log m)$.
Algorithm 4 Second-order kernel quantities from prefix–suffix row sums
1:Single-sample row sums $\{R_{i}^{\mathbf{X}},L_{i}^{\mathbf{X}}\}_{i=1}^{n}$, $\{R_{j}^{\mathbf{Y}},L_{j}^{\mathbf{Y}}\}_{j=1}^{m}$; Cross-sample accumulators $\{\mathcal{A}_{x\to y}(i),\mathcal{Z}_{x\to y}(i)\}_{i=1}^{n}$, $\{\mathcal{A}_{y\to x}(j),\mathcal{Z}_{y\to x}(j)\}_{j=1}^{m}$
2:Second-order kernel statistics: $\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}$, $\|\tilde{\mathbf{K}}_{XX}\mathbf{1}\|_{2}^{2}$, $(\mathbf{1}^{\top}\tilde{\mathbf{K}}_{XX}\mathbf{1})^{2}$, and analogues for $YY$ and XYXY
@@
[…]
20:return all above quantities
1:Samples $\mathbf{X}=\{X_{i}\}_{i=1}^{n}$, $\mathbf{Y}=\{Y_{j}\}_{j=1}^{m}$, kernel parameter $\sigma$
[…]
4:Sort $\mathbf{X}$ and $\mathbf{Y}$ to obtain $\mathbf{X}_{s}=\{x_{1}\leq\dots\leq x_{n}\}$ and $\mathbf{Y}_{s}=\{y_{1}\leq\dots\leq y_{m}\}$
@@
24:Compute $\mathbb{E}[g_{2,A}^{2}]$, $\mathbb{E}[g_{2,B}^{2}]$, $\mathbb{E}[g_{2,C}^{2}]$ using Eqs. (44)–(46) and Algorithm 4
@@
26:return $\widehat{\mathrm{MMD}}^{2}$, $\widehat{\operatorname{Var}}(\mathcal{T}_{1})+\widehat{\operatorname{Var}}(\mathcal{T}_{2})$, $\widehat{\operatorname{Var}}(\mathcal{T}_{1})$, $\widehat{\operatorname{Var}}(\mathcal{T}_{2})$
@@ 5.1 Variance analysis under heterogeneous settings
Table 1 presents the estimated variance under the equal sample size setting. The results confirm that our proposed statistic aligns perfectly with the classical unbiased MMD. However, a significant divergence appears at the variance level. The Sutherland estimator, while theoretically unbiased, relies on a fine grained decomposition of inner product terms, where each term is corrected independently. As shown in Table 1, this "term-wise" unbiasedness introduces severe numerical instability when the true variance is near zero. Specifically, for the Gaussian kernel at $n=15849$, the Sutherland estimator yields a non-physical negative value, a known artifact where bias-correction terms fluctuate beyond the magnitude of the raw estimate.
@@
$\widehat{\operatorname{Var}}$(MMD) (Gretton et al.) 2.448e-05 9.716e-07 3.933e-08 1.548e-09
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 1.179e-04 1.109e-06 2.624e-08 -1.233e-09
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 1.614e-04 2.983e-06 1.033e-07 1.847e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 2.280e-05 9.397e-07 3.853e-08 1.540e-09
@@
$\widehat{\operatorname{Var}}$(MMD) (Gretton et al.) 3.717e-05 1.485e-06 5.942e-08 2.345e-09
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 1.891e-04 1.828e-06 3.970e-08 -2.168e-09
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 2.543e-04 4.636e-06 1.557e-07 2.494e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 3.457e-05 1.408e-06 5.803e-08 2.331e-09
In contrast, our proposed method strictly follows the Hoeffding decomposition, separately estimating first- and second-order projection variances. This ensures positive semi-definiteness and superior numerical stability. Notably, our variance estimates are systematically larger than those of Gretton and Sutherland. This is because, in finite samples, empirical first-order projections retain substantial fluctuations due to cross-coupling effects among samples (Remark LABEL:remark1), whereas second-order components remain consistent with Gretton’s estimator.
@@ 5.2 Accuracy and efficiency of Laplacian-based accelerated euMMD
To validate the statistical precision of the proposed accelerated euMMD estimator, we conduct a comparative analysis under both $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ scenarios. We consider one-dimensional Laplace distributions where $\mathbb{P}\sim\mathrm{Laplace}(0,1)$ and $\mathbb{Q}\sim\mathrm{Laplace}(1,1)$, representing the alternative hypothesis with a location shift. To assess the performance under unbalanced conditions, the sample sizes are set such that $m=1.2n$, with $n\in\{10,100,1000,10000\}$.
Table 7 demonstrates that the proposed euMMD statistic is algebraically equivalent to the classical unbiased MMD, yielding identical values across all test cases. In terms of variance estimation, our accelerated framework provides a high-fidelity decomposition into first-order ($\mathcal{T}_{1}$) and second-order ($\mathcal{T}_{2}$) components.
A subtle numerical fluctuation is observed at an extremely small sample size ($n=10$), where the $\mathcal{T}_{2}$ component yields a slight negative value. This is attributed not to the theoretical properties of the U-statistic, but to numerical precision limits inherent in the accelerated computational path. Specifically, euMMD replaces the $O(n^{2})$ matrix summation with a sorting-based prefix sum accumulation. In the small sample regime, the floating-point rounding errors incurred during the sequential subtraction of sorted terms can lead to minor precision loss. However, as $n$ increases to 100 and beyond, these artifacts vanish entirely. The resulting variance estimates exhibit the expected $O(n^{-1})$ decay and achieve perfect alignment with classical results. This confirms that euMMD maintains rigorous statistical accuracy while significantly enhancing computational efficiency for large-scale applications.
Table 7: Comparison between MMD and euMMD
Sample size 10 100 1000 10000
MMD 5.206e-02 1.272e-01 1.193e-01 1.030e-01
euMMD 5.206e-02 1.272e-01 1.193e-01 1.030e-01
$\widehat{\operatorname{Var}}$(MMD) 1.465e-02 1.368e-03 1.202e-04 1.045e-05
$\widehat{\operatorname{Var}}$(euMMD) 9.790e-03 1.365e-03 1.202e-04 1.045e-05
$\widehat{\operatorname{Var}}$(MMD)($\mathcal{T}_{2}$) 3.355e-03 2.772e-05 2.740e-07 2.843e-09
$\widehat{\operatorname{Var}}$(euMMD)($\mathcal{T}_{2}$) -1.503e-03 2.444e-05 2.708e-07 2.840e-09

### v2

@@ 4.2.2 Accelerated computation of second-order components
$\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}$ $=\sum_{i=1}^{n}(R_{i}^{\mathrm{sq},\mathbf{X}}+L_{i}^{\mathrm{sq},\mathbf{X}}),$ (59)
@@
$\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}$ $=\sum_{j=1}^{m}(R_{j}^{\mathrm{sq},\mathbf{Y}}+L_{j}^{\mathrm{sq},\mathbf{Y}}),$ (62)
@@
$\|\mathbf{K}_{XY}\|_{F}^{2}$ $=\sum_{i=1}^{n}(\mathcal{A}^{\mathrm{sq}}_{x\to y}(i)+\mathcal{Z}^{\mathrm{sq}}_{x\to y}(i)),$ (65)
@@
Based on the theoretical analysis, Algorithm 4 presents an accelerated procedure for computing the Maximum Mean Discrepancy (MMD) with the Laplace kernel together with its full Hoeffding variance estimator, reducing the computational complexity from the naive $\mathcal{O}(N^{2})$ to $\mathcal{O}(n\log n+m\log m)$.
1:Samples $\mathbf{X}$, $\mathbf{Y}$, parameter $\sigma$
@@
4:Sort $\mathbf{X},\mathbf{Y}\to\mathbf{X}_{s},\mathbf{Y}_{s}$
@@
[…]
32:return $\widehat{\mathrm{MMD}}^{2},\widehat{\operatorname{Var}}(\mathcal{T}_{1})+\widehat{\operatorname{Var}}(\mathcal{T}_{2})$
@@ 5.1 Variance analysis under heterogeneous settings
Table 1 presents the estimated variance under the equal sample size setting. The results confirm that our proposed statistic aligns perfectly with the classical unbiased MMD. However, a significant divergence appears at the variance level. The Sutherland estimator, while theoretically unbiased, relies on a fine-grained decomposition of inner product terms, where each term is corrected independently. As shown in Table 1, this unbiasedness introduces severe numerical instability when the true variance is near zero. Specifically, for the Gaussian kernel at $n=3162$, the Sutherland estimator yields an abnormally low value ($6.170\times 10^{-10}$), a known artifact where bias-correction terms fluctuate beyond the magnitude of the raw estimate.
In contrast, our proposed method strictly follows the Hoeffding decomposition, separately estimating first- and second-order projection variances. This ensures positive semi-definiteness and superior numerical stability. Notably, our variance estimates are systematically larger than those of Gretton and Sutherland. This is because, in finite samples, empirical first-order projections retain substantial fluctuations due to cross-coupling effects among samples (Remark 1), whereas second-order components remain consistent with Gretton’s estimator.
@@
$\widehat{\operatorname{Var}}$(MMD) (Gretton et al.) 2.145e-05 8.782e-07 3.725e-08 1.499e-09
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 3.225e-05 1.324e-06 1.154e-08 2.917e-09
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 8.456e-05 3.331e-06 8.814e-08 5.970e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 2.664e-05 1.007e-06 3.831e-08 1.527e-09
@@
$\widehat{\operatorname{Var}}$(MMD) (Gretton et al.) 3.363e-05 1.337e-06 5.647e-08 2.254e-09
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 4.652e-05 1.622e-06 6.170e-10 2.862e-09
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 1.296e-04 4.761e-06 1.160e-07 7.457e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 4.232e-05 1.574e-06 5.771e-08 2.297e-09
@@ 5.2 Accuracy and efficiency of Laplacian-based accelerated euMMD
To validate the statistical precision of the proposed accelerated euMMD estimator, we conduct a comparative analysis under the alternative hypothesis $\mathcal{H}_{1}$. We consider one-dimensional Laplace distributions where $\mathbb{P}\sim\mathrm{Laplace}(0,1)$ and $\mathbb{Q}\sim\mathrm{Laplace}(1,1)$, representing a location shift. To assess the performance under imbalanced conditions, the sample sizes are set at a ratio of $m=1.2n$, with $n\in\{10,100,1000,10000\}$.
As shown in Table 7, the proposed euMMD statistic and its variance components ($\mathcal{T}_{1}$ and $\mathcal{T}_{2}$) are algebraically identical to the classical matrix-form unbiased MMD estimator. This alignment confirms that our $O(N\log N)$ framework is not an approximation but an exact acceleration of the finite-sample U-statistic.
Notably, in the small sample regime ($n=10$), the estimator maintains rigorous consistency. By carefully aligning the prefix-suffix accumulators to exclude diagonal self-similarity terms $k(x_{i},x_{i})$, the resulting second-order component $\mathcal{T}_{2}$ remains positive and numerically precise. As $n$ increases, the variance estimates exhibit the expected $O(n^{-1})$ decay, with the first-order component $\mathcal{T}_{1}$ dominating the total variance. These results demonstrate that euMMD preserves the full statistical fidelity of the MMD while significantly enhancing computational efficiency for large-scale data analysis.
Table 7: Numerical Consistency Verification: Matrix-form MMD vs. Accelerated euMMD
Sample size ($n,m$) (10, 12) (100, 120) (1000, 1200) (10000, 12000)
$\widehat{\mathrm{MMD}}^{2}$ (Matrix) 5.206e-02 1.272e-01 1.193e-01 1.030e-01
$\widehat{\mathrm{MMD}}^{2}$ (euMMD) 5.206e-02 1.272e-01 1.193e-01 1.030e-01
$\widehat{\operatorname{Var}}_{\text{total}}$ (Matrix) 1.465e-02 1.368e-03 1.202e-04 1.045e-05
$\widehat{\operatorname{Var}}_{\text{total}}$ (euMMD) 1.465e-02 1.368e-03 1.202e-04 1.045e-05
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Matrix) 3.355e-03 2.772e-05 2.740e-07 2.843e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (euMMD) 3.355e-03 2.772e-05 2.740e-07 2.843e-09
Note: All results are identical up to machine precision, validating the exactness of the $O(N\log N)$ implementation.